# Gradient Minimization Using Higher Order MethodsYao Ji and Guanghui Lan

### v1

@@ 4 Parameter-free algorithm for convex problems
    Unknown initial optimality gap $D$: However, in general $\text{dist}(x_{0},X^{*})$ is unknown, and we may use $D$ to estimate it. If $D>\text{dist}(x_{0},X^{*}),$ by choosing $\sigma_{1}={\varepsilon}/[3(9D)^{p+\nu-1}],$ we can compute an approximate solution $\widehat{x}$ using Algorithm 4 such that $\|\nabla f(\widehat{x})\|\leq\varepsilon$ after
    $\mathcal{O}\left(\left[\tfrac{\max\{{pL_{p,\nu}}(f),L_{0},\theta\}[C_{p}D^{p+\nu-1}]}{\varepsilon}\right]^{\frac{1}{p+\nu}}\right)$
    calls to the $p$-th order oracle, which is worse than (46). Next we show that we can avoid such case by choosing $D$ properly. Specifically, define $D^{p+\nu-1}\coloneqq\|\nabla f_{1}({x}_{1}^{2})\|/L_{1,2},$ by (36) with $s=1,N_{1}=1,$ we have
@@
    where $x^{*}$ is any optimal solution in $X^{*}.$ Therefore, we have $D\leq\text{dist}(x_{0},X^{*}).$ If $D<\text{dist}(x_{0},X^{*}),$ by choosing $\sigma_{1}={\varepsilon}/[3(9D)^{p+\nu-1}],$ the computed approximate solution $\widehat{x}$ only satisfies $\|\nabla f(\widehat{x})\|\leq\varepsilon[\text{dist}(x_{0},X^{*})/{D}]^{p+\nu-1}.$ Therefore, it does not attain the desired accuracy $\varepsilon.$ In this case, we simply discard all previous results and restart the computation with a larger guess $D,$ such a procedure that gradually increases $D$ is called a “guess-and-check” procedure. It is first used parameter-free optimal gradient minimization using first order method [21]. We will show below that such a strategy maintains the same oracle complexity as if $\operatorname{dist}(x_{0},X^{*})$ were known, while still guaranteeing $\|\nabla f(\widehat{x})\|\leq\varepsilon$.
@@
Suppose the Assumptions in Theorem 4.1. Then Algorithm 5 computes a solution $\widehat{x}$ such that $\|\nabla f(\widehat{x})\|\leq\varepsilon$ within at most (46) calls to the $p$-th order oracle.
@@
Observe that after at most $T$ calls to Algorithm 4, Algorithm 5 will terminate, where $T$ satisfies $D_{T}=4^{T}D_{0}\geq\text{dist}(x_{0},X^{*}),$ and $D_{T-1}=4^{T-1}D_{0}\leq\text{dist}(x_{0},X^{*}),$ therefore, we have $T-1\leq\lceil\log_{4}\tfrac{\text{dist}(x_{0},X^{*})}{D_{0}}\rceil.$ By Theorem 4.1, in total, Algorithm 5 requires
$c\textstyle\sum_{t=1}^{T-1}\left[\tfrac{c_{\mathcal{A}}\max\{pL_{p,\nu}(f),L_{0},\theta\}(p+\nu)[3(9D_{0}\cdot 4^{t})^{p+\nu-1}]}{4{\varepsilon}}\right]^{\frac{1}{p+\nu}}$
$\leq c\left[\tfrac{c_{\mathcal{A}}\max\{pL_{p,\nu}(f),L_{0},\theta\}(p+\nu)[3(9D_{0})^{p+\nu-1}]}{4{\varepsilon}}\right]^{\frac{1}{p+\nu}}\tfrac{4^{\frac{(T-1)(p+\nu-1)}{p+\nu}}-1}{4^{\frac{p+\nu-1}{p+\nu}}-1}$
$\leq c\left[\tfrac{c_{\mathcal{A}}\max\{pL_{p,\nu}(f),L_{0},\theta\}(p+\nu)[3(9\text{dist}(x_{0},X^{*}))^{p+\nu-1}]}{4{\varepsilon}}\right]^{\frac{1}{p+\nu}}\tfrac{1}{4^{\frac{p+\nu-1}{p+\nu}}-1},$
calls to the $p$-th order oracle, where $c$ is a universal constant. This concludes the proof.
@@ 5 Gradient Minimization Framework for Uniformly Convex Problems
In this section, we show that by using optimal (sub)gradient norm convergence results of the high-order AR framework for convex composite problems, we can obtain new fast (sub)gradient norm convergence rates for composite problems satisfying a certain regularity condition.
5.1 Gradient minimization for convex composite problems
Consider a class of composite problems (3) and in addition, $f$ is uniformly convex on $\mathbb{R}^{n}$ of degree $q$ with parameter $\sigma_{q}(f)>0$ [31, 4, 33, 35], i.e.,
(47) $f(x)\geq f(y)+\langle g_{y},x-y\rangle+\tfrac{\sigma_{q}(f)}{q}\|x-y\|^{q},\quad\forall\,x,y\in\mathbb{R}^{n},g_{y}\in\partial f(y).$
Consider restarting the high-order AR framework Algorithm 3 whenever the (sub)gradient norm is reduced by half. Specifically, in each epoch $k,$ we run $m_{k}$ iterations of high-order AR framework Algorithm 3 with the initialization $x_{k-1},$ i.e., the output of the previous epoch k1.k-1. By a careful choice of the epoch length $m_{k}$, which depends on the relation between the upper smoothness level $p$ and the lower uniform convexity level $q$, we can ensure the existence of a subgradient sequence with exponentially decaying norms. We present below the high-order AR framework for uniformly convex composite problems together with its convergence guarantee.
Algorithm 6 Uniformly convex accumulative regularization
1: Initial point $x_{0}\in\mathbb{R}^{n},$ $L_{p+1}(g),\sigma_{q}(f).$
@@ 5.1 Gradient minimization for convex composite problems
3:   Compute an approximate solution $(x_{k},\nu_{k})$ of $f$ with initialization $(x_{k-1},\nu_{k-1})$ by running high-order AR framework in Algorithm 3 for $m_{k}$ iterations, where $m_{k}=\max\left\{\left(\tfrac{qL_{p+1}(g)}{{\sigma_{q}(f)}}\right)^{\frac{2}{3p+1}}\left({\tfrac{q\|\nu_{k-1}\|}{2\sigma_{q}(f)}}\right)^{\frac{2(p-q+1)}{(3p+1)(q-1)}},1\right\}.$
@@
5: output $\widehat{x}$
@@
Suppose that $\|\nu_{0}\|\geq\varepsilon.$ Then, Algorithm 6 can compute a point $\widehat{x}$ such that $\|\widehat{\nu}\|\leq\varepsilon$ within at most
@@
calls to the $p$-th order oracle if $q=p+1;$ and at most
@@
calls to the $p$-th order oracle if $q>p+1;$ and at most
@@
calls to the $p$-th order oracle if $q<p+1.$
@@
By adding two copies of inequality (47), we have
(48) $\langle x-y,g_{x}-g_{y}\rangle\geq\tfrac{2\sigma_{q}(f)}{q}\|x-y\|^{q},\quad\forall\,\,x,y\in\mathbb{R}^{n},$
and for all $g_{x}\in\partial f(x),g_{y}\in\partial f(y).$ Substituting $x=x_{k-1}$ and $y=x^{*}$ into (48), and using the Cauchy-Swartcz inequality and optimality condition $0\in\partial f(x^{*}),$ we can choose $g_{x^{*}}=0$ and thus, we have
(49) $\|x_{k-1}-x^{*}\|\leq\left({\tfrac{q\|g_{k-1}\|}{2\sigma_{q}(f)}}\right)^{\frac{1}{q-1}},$
where $g_{k-1}$ is an arbitrary sub-gradient of $f$ at point $x_{k-1}.$ Let us prove by induction that there exists a sequence $\{\nu_{k}\}_{k\geq 1}$ such that $\nu_{k}\in\partial f(x_{k})$ and $\|\nu_{k}\|\leq\|\nu_{0}\|/2^{k}.$ Suppose it is true for k1.k-1. At epoch $k,$ we restart the high-order AR framework Algorithm 3 with the initial point $x_{k-1}.$ By Theorem 3.11, Algorithm 3 can compute an approximate solution $x_{k}$ together with $\nu_{k}$ such that
@@
Given that $g_{k-1}$ can be any sub-gradient of $f$ at point $x_{k-1},$ then we choose $g_{k-1}=\nu_{k-1}.$ By the choice of $m_{k}$ in Algorithm 6, we have $\|\nu_{k}\|\leq{{\|\nu_{k-1}\|}/{2}}\leq\|\nu_{0}\|/2^{k}.$ By simply restarting the high-order AR framework whenever the norm of the approximate solution $\nu_{k}$ is reduced by half, we have the following convergence guarantees.
Case I: $q=p+1.$ In this case, $m_{k}\equiv m_{0}=\left[{qL_{p+1}(g)}/{{\sigma_{p+1}(f)}}\right]^{\frac{2}{3p+1}},$ thus the convergence rate is linear. We can obtain an approximate solution $\widehat{x}$ such that $\|\hat{\nu}\|\leq\varepsilon$ within at most $\mathcal{O}(\log{\|\nu_{0}\|}/{\varepsilon})$ epochs.
Case II: $q>p+1.$ In this case, $m_{k}$ increases exponentially, to reach a point ${x}_{K}$ such that $\|\nu_{K}\|\leq\varepsilon,$ we need
$\textstyle\sum_{k=1}^{K}m_{k}$ $=\left(\tfrac{qL_{p+1}(g)}{{\sigma_{q}(f)}}\right)^{\frac{2}{3p+1}}\left(\tfrac{\sigma_{q}(f)}{q\|\nu_{0}\|}\right)^{\frac{2(q-p-1)}{(3p+1)(q-1)}}\textstyle\sum_{k=1}^{K}2^{\frac{2(q-p-1)k}{(3p+1)(q-1)}}$
@@
calls to the $p$-th order oracle, where in (a), we used $\|\nu_{K}\|\leq\|\nu_{0}\|/2^{K}\leq\varepsilon.$
Case III: $q<p+1.$ In this case, $m_{k}$ decays exponentially until it reaches $1,$ then, the algorithm converges superlinearly. Thus, the algorithm has two phases. Let $k=K_{1}$ denote the final iteration count of Phase I, where $m_{K_{1}}=1$ and $m_{K_{1}-1}\geq 1,$ we have
(50) $\left(\tfrac{qL_{p+1}(g)}{{\sigma_{q}(f)}}\right)^{\frac{2}{3p+1}}\left(\tfrac{q\|\nu_{K_{1}-1}\|}{2\sigma_{q}(f)}\right)^{{\frac{2(p-q+1)}{(3p+1)(q-1)}}}\leq 1,\quad\text{and}\quad m_{K_{1}}=1.$
At epoch $K_{1},$ by Theorem 3.11, we have
$\|\nu_{K_{1}}\|$ $\,\,\,\,\leq\tfrac{L_{p+1}(g)\|x_{K_{1}-1}-x^{*}\|^{p}}{m_{K_{1}}^{\frac{3p+1}{2}}}={L_{p+1}(g)\|x_{K_{1}-1}-x^{*}\|^{p}}$
@@
Define $H=\left(\tfrac{1}{L_{p+1}(g)}\right)^{\frac{q-1}{p+1-q}}\left(\tfrac{2\sigma_{q}(f)}{q}\right)^{\frac{p}{p+1-q}},$ then we have $\tfrac{\|\nu_{K_{1}}\|}{H}\leq\tfrac{1}{2}.$ The total iterations of Phase I reads
$\textstyle\sum_{k=1}^{K_{1}}m_{k}$ $\leq\textstyle\sum_{k=1}^{K_{1}}\left(\tfrac{qL_{p+1}(g)}{{\sigma_{q}(f)}}\right)^{\frac{2}{3p+1}}\left(\tfrac{q\|\nu_{0}\|}{2^{k}\sigma_{q}(f)}\right)^{{\frac{2(p-q+1)}{(3p+1)(q-1)}}}$
@@
Then, Algorithm 6 enters into Phase II: superlinear convergence phase. For all $k\geq K_{1},$ $m_{k}=1,$ by Theorem 3.11, $\|\nu_{k}\|\leq L_{p+1}(g)\left({{q\|\nu_{k-1}\|}/{2\sigma_{q}(f)}}\right)^{\frac{p}{q-1}},$ thus ${\|\nu_{k}\|}/{H}\leq\left({\|\nu_{k-1}\|}/{H}\right)^{\frac{p}{q-1}}.$ Therefore, for all $k\geq K_{1},$ we have $\|\nu_{k}\|\leq H\left(\tfrac{1}{2}\right)^{(\frac{p}{q-1})^{k-K_{1}}}.$ For any $\varepsilon>0,$ to reach a point $\widehat{x}$ to satisfy $\|\widehat{\nu}\|\leq\varepsilon,$ we need $k\geq\tfrac{\log\log\tfrac{H}{\varepsilon}-\log\log 2}{\log{\left(\frac{p}{q-1}\right)}}+K_{1}$ iterations.
Depending on the relation between $p+1$ and $q,$ Algorithm 6 can exhibit linear or superlinear convergence rate. Such complexities in terms of the gradient norm have not been derived before.
5.2 parameter-free gradient minimization for uniformly convex function
In this section, we develop parameter-free and inexact method for uniformly convex function. In particular, we do not assume the knowledge of Lipschitz continuous parameter and uniform convex parameter. We will utilize the parameter-free gradient minimization methods developed from Section 4, and use a guess and search procedure to estimate the uniform convex parameter.
For simplicity, we will focus on the unconstrained problem (1) where $f$ is $p$-times differentiable with a Lipschitz continuous $p$-th derivative (2), and furthermore $f$ is uniformly convex with order $q=p+1.$ We have the following parameter-free and inexact algorithm for uniformly convex function.
Algorithm 7 A parameter-free uniformly convex accumulative regularization
1: Initial point $f,\varepsilon,x_{0},\sigma_{p+1,0},L_{0}.$

### v2

@@ 4 Parameter-free algorithm for convex problems
    Unknown initial optimality gap $D$: However, in general $\operatorname{dist}(x_{0},X^{*})$ is unknown, and we may use $D$ to estimate it. If $D>\operatorname{dist}(x_{0},X^{*}),$ by choosing $\sigma_{1}={\varepsilon}/[3(9D)^{p+\nu-1}],$ we can compute an approximate solution $\widehat{x}$ using Algorithm 4 such that $\|\nabla f(\widehat{x})\|\leq\varepsilon$ after
    $\mathcal{O}\left(\left[\tfrac{\max\{{pL_{p,\nu}}(f),L_{0},\theta\}D^{p+\nu-1}}{\varepsilon}\right]^{\frac{1}{p+\nu}}\right)$
    calls to the $p$-th-order oracle, which is worse than (46). Next we show that we can avoid such case by choosing $D$ properly. Specifically, run $\mathcal{A}_{1}$ with $N_{1}=1,$ set $D^{p+\nu-1}\coloneqq\|\nabla f_{1}({x}_{1}^{2})\|/L_{1,2},$ by (36), we have
@@
    where $x_{1}^{2}$ is the computed approximate solution at epoch $1$ after $2$ iterations, and $x^{*}$ is any optimal solution in $X^{*}.$ Therefore, we have $D\leq\operatorname{dist}(x_{0},X^{*}).$ If $D<\operatorname{dist}(x_{0},X^{*}),$ by choosing $\sigma_{1}={\varepsilon}/[3(9D)^{p+\nu-1}],$ the computed approximate solution $\widehat{x}$ only satisfies $\|\nabla f(\widehat{x})\|\leq\varepsilon[\operatorname{dist}(x_{0},X^{*})/{D}]^{p+\nu-1}.$ Therefore, it does not attain the desired accuracy $\varepsilon.$ In this case, we simply discard all previous results and restart the computation with a larger guess $D,$ this increasing $D$ schedule is a “guess-and-check” procedure. It was first used for parameter-free optimal gradient minimization with first-order methods [lan2024optimalparameterfreegradientminimization]. We will show below that such a strategy maintains the same oracle complexity as if $\operatorname{dist}(x_{0},X^{*})$ were known, while still guaranteeing $\|\nabla f(\widehat{x})\|\leq\varepsilon$.
@@
Suppose the assumptions in Theorem 4.1. Then Algorithm 5 computes a solution $\widehat{x}$ such that $\|\nabla f(\widehat{x})\|\leq\varepsilon$ within at most (46) calls to the $p$-th-order oracle.
@@
Observe that after at most $T$ calls to Algorithm 4, Algorithm 5 will terminate, where $T$ satisfies $D_{T}=4^{T}D_{0}\geq\operatorname{dist}(x_{0},X^{*}),$ and $D_{T-1}=4^{T-1}D_{0}\leq\operatorname{dist}(x_{0},X^{*}),$ therefore, we have $T-1\leq\lceil\log_{4}{\operatorname{dist}(x_{0},X^{*})}/{D_{0}}\rceil.$ By Theorem 4.1, in total, Algorithm 5 requires
$c\textstyle\sum_{t=1}^{T}\left[\tfrac{c_{\mathcal{A}}\max\{pL_{p,\nu}(f),L_{0},\theta\}(p+\nu)[3(9D_{0}\cdot 4^{t})^{p+\nu-1}]}{4{\varepsilon}}\right]^{\frac{1}{p+\nu}}$
$\leq c\left[\tfrac{c_{\mathcal{A}}\max\{pL_{p,\nu}(f),L_{0},\theta\}(p+\nu)[3(9D_{0})^{p+\nu-1}]}{4{\varepsilon}}\right]^{\frac{1}{p+\nu}}\tfrac{4^{{T(p+\nu-1)}/{(p+\nu)}}\cdot 4^{{(p+\nu-1)}/{(p+\nu)}}}{4^{{(p+\nu-1)}/{(p+\nu)}}-1}$
$\leq c\left[\tfrac{c_{\mathcal{A}}\max\{pL_{p,\nu}(f),L_{0},\theta\}(p+\nu)[3(9\operatorname{dist}(x_{0},X^{*}))^{p+\nu-1}]}{4{\varepsilon}}\right]^{\frac{1}{p+\nu}}\tfrac{16^{{(p+\nu-1)}/{(p+\nu)}}}{4^{{(p+\nu-1)}/{(p+\nu)}}-1}$
calls to the $p$-th-order oracle, where $c$ is a universal constant.
@@ 5 Gradient Minimization Framework for Uniformly Convex Problems
In this section, leveraging the (sub)gradient norm convergence results of the high-order AR algorithms established in Section 3 and Section 4 for convex composite problems, we obtain fast (sub)gradient-norm rates for composite problems that satisfy a suitable regularity condition.
5.1 High-order AR for uniformly convex composite problems
Consider a class of composite problems (3), and in addition, $f$ is uniformly convex on $\mathbb{R}^{n}$ of degree $q$ with parameter $\sigma_{q}(f)>0$ [polyak1966existence, aze1995uniformly, vladimirov1978uniformly, zǎlinescu1983uniformly], i.e.,
(47) $f(x)\geq f(y)+\langle g_{y},x-y\rangle+\tfrac{\sigma_{q}(f)}{q}\|x-y\|^{q},\quad\forall\,x,y\in\mathbb{R}^{n},\quad\forall\,g_{y}\in\partial f(y),$
where $q\geq 2.$ Here, we adopt the characterization of uniform convexity via the subgradient inequality; see for example [bauschke2020correction, Proposition 17.26].
Algorithm 6 High-order AR for uniformly convex functions
1: Initial point $x_{0}\in\mathbb{R}^{n}$ and initial subgradient $\nu_{0}\in\partial f(x_{0});$ $L_{p+1}(g);\sigma_{q}(f).$
@@ 5.1 High-order AR for uniformly convex composite problems
3:   Compute an approximate solution $(x_{k},\nu_{k})$ of $f$ with initialization $(x_{k-1},\nu_{k-1})$ by running high-order AR framework Algorithm 3 for $m_{k}$ iterations, where $m_{k}=\max\left\{\left(\tfrac{qL_{p+1}(g)}{{\sigma_{q}(f)}}\right)^{\frac{2}{3p+1}}\left({\tfrac{q\|\nu_{k-1}\|}{2\sigma_{q}(f)}}\right)^{\frac{2(p-q+1)}{(3p+1)(q-1)}},1\right\}.$
@@
5: return $\widehat{x}$
To solve this problem, we employ a restart scheme: restart the high-order AR framework Algorithm 3 whenever the (sub)gradient norm decreases by half. Specifically, we consider Algorithm 6, where for each epoch $k,$ we run $m_{k}$ iterations of the high-order AR framework Algorithm 3 with the initialization $x_{k-1},$ i.e., the output of the previous epoch, k1.k-1. With a suitable selection of the epoch lengths $m_{k}$, determined by the upper smoothness order $p$ and the lower uniform convexity order $q$, we obtain a sequence of subgradients with exponentially decaying norms across epochs. We next present the convergence guarantee for Algorithm 6.
@@
[…]
@@
calls to the $p$-th-order oracle if $q=p+1;$ and at most
@@
calls to the $p$-th-order oracle if $q>p+1;$ and at most
@@
calls to the $p$-th-order oracle if $q<p+1.$
@@
By adding two copies of the inequality (47), we have
(48) $\langle x-y,g_{x}-g_{y}\rangle\geq\tfrac{2\sigma_{q}(f)}{q}\|x-y\|^{q},\quad\forall\,\,x,y\in\mathbb{R}^{n},\quad\forall\,g_{x}\in\partial f(x),\,g_{y}\in\partial f(y).$
Substituting $x=x_{k-1}$ and $y=x^{*}$ into (48), and using the Cauchy-Schwarz inequality and the optimality condition $0\in\partial f(x^{*}),$ we can choose $g_{x^{*}}=0$ and thus, we have
(49) $\|x_{k-1}-x^{*}\|\leq\left({\tfrac{q\|g_{k-1}\|}{2\sigma_{q}(f)}}\right)^{\frac{1}{q-1}},\quad\forall\,g_{k-1}\in\partial f(x_{k-1}).$
Let us prove by induction that there exists a sequence $\{\nu_{k}\}_{k\geq 1}$ such that $\nu_{k}\in\partial f(x_{k})$ and $\|\nu_{k}\|\leq\|\nu_{0}\|/2^{k}.$ Suppose it holds for k1.k-1. At epoch $k,$ we restart Algorithm 3 with the initial point $x_{k-1}.$ By Theorem 3.11, Algorithm 3 can compute an approximate solution $x_{k}$ together with $\nu_{k}\in\partial f(x_{k})$ such that
@@
Given that $g_{k-1}$ can be any subgradient of $f$ at the point $x_{k-1},$ we choose $g_{k-1}=\nu_{k-1}.$ By the choice of $m_{k}$ in Algorithm 6, we have $\|\nu_{k}\|\leq{{\|\nu_{k-1}\|}/{2}}\leq\|\nu_{0}\|/2^{k}.$ We next bound the iteration complexity.
Case I: $q=p+1.$ In this case, $m_{k}\equiv m_{0}=\left[{qL_{p+1}(g)}/{{\sigma_{p+1}(f)}}\right]^{\frac{2}{3p+1}},$ so the convergence rate is linear. We can obtain an approximate solution $\widehat{x}$ such that $\|\widehat{\nu}\|\leq\varepsilon$ within at most $\mathcal{O}(\log{\|\nu_{0}\|}/{\varepsilon})$ epochs.
Case II: $q>p+1.$ In this case, $m_{k}$ increases exponentially, and to reach a point ${x}_{K}$ such that $\|\nu_{K}\|\leq\varepsilon,$ we need
$\textstyle\sum_{k=1}^{K}m_{k}$ $=\left(\tfrac{qL_{p+1}(g)}{{\sigma_{q}(f)}}\right)^{\frac{2}{3p+1}}\left(\tfrac{\sigma_{q}(f)}{q\|\nu_{0}\|}\right)^{\frac{2(q-p-1)}{(3p+1)(q-1)}}\textstyle\sum_{k=1}^{K}2^{\frac{2(q-p-1)(k-1)}{(3p+1)(q-1)}}$
@@
calls to the $p$-th-order oracle, where in (a), we used $\varepsilon<\|\nu_{K-1}\|\leq\|\nu_{0}\|/2^{K-1}.$
Case III: $q<p+1.$ In this case, $m_{k}$ decays exponentially until it reaches $1,$ then, the algorithm converges superlinearly. Thus, the algorithm has two phases. Let $k=K_{1}$ denote the final iteration count of Phase I, where $m_{K_{1}}=1$ and $m_{K_{1}-1}>1,$ we have
(50) $m_{K_{1}}=\max\left\{\left(\tfrac{qL_{p+1}(g)}{{\sigma_{q}(f)}}\right)^{\frac{2}{3p+1}}\left(\tfrac{q\|\nu_{K_{1}-1}\|}{2\sigma_{q}(f)}\right)^{{\frac{2(p-q+1)}{(3p+1)(q-1)}}},1\right\}=1.$
Then, at epoch $K_{1},$ by Theorem 3.11, we have
$\|\nu_{K_{1}}\|$ $\,\,\,\,\leq{L_{p+1}(g)\|x_{K_{1}-1}-x^{*}\|^{p}}$
@@
The total number of iterations in Phase I is
$\textstyle\sum_{k=1}^{K_{1}}m_{k}$ $\leq\textstyle\sum_{k=1}^{K_{1}-1}\left(\tfrac{qL_{p+1}(g)}{{\sigma_{q}(f)}}\right)^{\frac{2}{3p+1}}\left(\tfrac{q\|\nu_{0}\|}{2^{k-1}\sigma_{q}(f)}\right)^{{\frac{2(p-q+1)}{(3p+1)(q-1)}}}+1$
@@
Next, Algorithm 6 enters Phase II: the superlinear convergence phase. For all $k\geq K_{1},$ $m_{k}=1,$ furthermore, by Theorem 3.11 and (49), we have
$\|\nu_{k}\|\leq L_{p+1}(g)\left(\tfrac{q\|\nu_{k-1}\|}{2\sigma_{q}(f)}\right)^{\frac{p}{q-1}}.$
Define $H\coloneqq{L_{p+1}(g)}^{\frac{q-1}{q-(p+1)}}\left(\tfrac{q}{2\sigma_{q}(f)}\right)^{\frac{p}{q-(p+1)}},$ thus, we have
[…]
Therefore, for all epochs $k\geq K_{1},$ we have $\|\nu_{k}\|\leq H2^{-(\frac{p}{q-1})^{k-K_{1}}}$ and $m_{k}=1.$ For any $\varepsilon>0,$ to reach a point $\widehat{x}$ that satisfies $\|\widehat{\nu}\|\leq\varepsilon,$ we need
$t\geq\tfrac{1}{\log{\left(\frac{p}{q-1}\right)}}\left({\log\log\tfrac{H}{\varepsilon}-\log\log 2}\right){}+\textstyle\sum_{k=1}^{K_{1}}m_{k}$
iterations. This concludes the proof.
Depending on the relation between $p+1$ and $q,$ Algorithm 6 can exhibit linear, sublinear or superlinear convergence rates. Such complexities in terms of the gradient norm have not been derived before.
5.2 Parameter-free gradient minimization for uniformly convex functions
In this section, we develop a parameter-free, inexact method for uniformly convex problems that does not assume knowledge of the Lipschitz constant or the uniform convexity parameter. We build on Section 4 and use a guess-and-check scheme to estimate the uniform convexity parameter.
For simplicity, we will focus on the unconstrained problem (1) where $f$ is $p$-times differentiable with a Lipschitz continuous $p$-th derivative (2), and assume that $f$ is uniformly convex with order $q=p+1.$ Below we present Algorithm 7, a parameter-free and inexact algorithm tailored to uniformly convex functions, together with its convergence guarantees.
Algorithm 7 Parameter-free uniformly convex accumulative regularization
1: Initial point $x_{0};$ initial uniform convexity guess $\sigma_{p+1,0};$ initial line-search value $L_{0};$ target accuracy $\varepsilon.$